-t points osUniphyst 5.2 P022 My Nese Two forces ofh- Two forces of Fi--N and F2- o-nr. and?2-110.00-JNact on an obje t. Fi d the third forcei, j) N act on an object. Find the third force F3 (in N) that is needed to balance the first two forces. (Express your answer in vector form) nN) that eeded to balance the rst two forces (Express your answer nvector forn. Additional Materials eBook Suomit Answsr Save Piogress Practice Another Version My Nate Suppose two chldren push horizontally, but in exactly opposite directions, on a third child in a wagon. The first child exerts a force of 70.0 N, the second exerts a force of 86.0 N, friction is 10.0 N, and the mass of the third child plus wagon is 23.0 kg ta) What is the system of interest if the acceleration of the child in the wapon is to be calculated? (Select all that apply) the children outside the wagon the child in the wagon the wagon (b) Calculate the acceleration (in m/s). (Enter the magnitude.) (c) What would the acceleration be (in m/s2) if friction were 16.0 N? (Enter the magnitude.) m/s
Answer #1

1.

We need Force F3 such that balance Forces F1 and F2, So that means

F1 + F2 + F3 = 0

F3 = -(F1 + F2)

Given that

F1 = (74.0/sqrt 2)*(i - j)

F2 = (110.0/sqrt 2)*(i - j)

So,

F1 + F2 = (74.0/sqrt 2)*(i - j) + (110.0/sqrt 2)*(i - j)

F1 + F2 = (74.0/sqrt 2 + 110.0/sqrt 2) i - (74.0/sqrt 2 + 110.0/sqrt 2) j

F1 + F2 = (184.0/sqrt 2) i - (184.0/sqrt 2) j

F1 + F2 = (184.0/sqrt 2)*(i - j)

Now,

F3 = -(F1 + F2)

F3 = -[(184.0/sqrt 2)*(i - j)]

F3 = (184.0/sqrt 2)*(-i + j)

So in vector form

184.0